Prairienet Banner

Implementation Plan

Thursday, 10-26-06

Accomplished:

Boot and nuked six computers: 9, 8, 7, 5, 4 and 2.
Cleaned four computers: 9, 8, 2, and 1.

Goals for the rest of this week:

Willy, call or e-mail to confirm dates, get answers to remaining questions for Site Survey, and confirm that the website was sent to the right address.

Goals for next week:

Wednesday (11-1-06):
Clean remaining computers (7, 6, 5, 4, and 3).
Boot and nuke remaining computers (6, 3, and 1).

Thursday (11-2-06):

Start fdisk

Wednesday, 11-1-06

Accomplished:

Cleaned three computers: 4, 5, and 7
Boot and nuked two computers: 6 and 3
We tried to boot and nuke the last computer (1); its official name is now “Demon Computer.” It wouldn’t boot and nuke. We tried different discs (floppy and cd), replaced cd-rom and floppy drives, and changed the boot order. We tried to see if the drivers were there but couldn’t get past the former owner’s password. We switched the jumper and deleted the old passwords twice; it didn’t work on either occasion. The plan: ask Martin for advice tomorrow in lab.

Thursday (11-2-06):

Clean computer 6 and put in cd-rom drive.
Fix Demon Computer (1).
fdisk and upload Windows 98 on as many computers as possible. (Goal: 6)

Thursday, 11-2-06

Accomplished:

Cleaned and added cd-rom to 6
Fdisked and put windows 98 on seven computers. Computers 1 and 2 (Khan and Sarek) are having problems. We replaced the hard drives.

Wednesday (11-8-06):

Get Khan and Sarek nuked.
Start updates.

Thursday (11-9-06):

Fdisk and Win98 Khan and Sarek.
Updates and drivers. Finish as many as possible.

Wednesday, 11-8-06

Accomplished:

Nothing. Despite changing out the hard drives, Khan and Sarek are still having the exact same problems. We were unable to nuke or format them. (This suggests the problem is not with the hard drive.)

Thursday (11-9-06):

If we can figure out what’s wrong with Khan and Sarek, great. If not, leave them alone until we can ask Martin for help.
Updates and drivers. Finish as many as possible.

Thursday, 11-9-06

Accomplished:

? Got Sarek working.
? Sarek formatted, win 98 installed, video, updates.
? Video and Windows updates on all computers but Khan.

Wednesday (11-15-06):

? Finish updates.
? Install software.
? Sound?

Thursday (11-16-06):

? Work on Khan.
? Ask Martin about Q’s .vdx file. What is it? Why do we keep getting that message?
? Finish M. Software. Start A-O Software.

Friday, 11-10-06 through Sunday, 11-12-06

Accomplished:

Created five posters to hang in the computer lab: Science and Math, English, Literature, History, and E-mail. They list websites that will help students with their Science, Math and English skills, find classic works of literature in a variety of languages in the public domain, and help them set up e-mail accounts.

Created a final software list and downloaded the software.

Tuesday, 11-14-06

Accomplished:

Burned four software discs for the computers. Here is the list with description.

Wednesday, 11-15-06

Question: Bobbi, would you have time tonight to get all possible sound drivers on a disc tonight? (I’ll also try to remember to bring my computer and some blank discs tomorrow.)

Accomplished:

Finished running updates on all computers but Khan. Put Fall 06 software (M. Software) on five computers. We couldn’t tell if the anti-virus software was actually downloading correctly so we still need to check that on all computers. (We put everything but anti-virus software on the five computers we did today.)

Thursday (11-16-06):

1. Upload the USB support disc on all computers. (It doesn’t come with Win. 98. If we want it on the computers, which we do, we’re going to have to download it from the floppy or cd-rom that’s floating around the lab.)
2. We’re going to try switching out the cables in Khan and running Boot and Nuke with a bootable Windows 98 cd tomorrow. (Ruth’s suggestion.) Maybe the computer just can’t run Linux. Otherwise, ask Martin for help.
3. Ask Martin about Q’s weird error messages. Ruth said they had that error message in E. St. Louis. She thinks we may have to download some software/drivers/something along those lines.
4. Get M. Software on the remaining four computers. (Update table accordingly.)
5. Get the anti-virus software on all computers.
6. Start putting A-O software on computers.
7. Sound drivers?!

Thursday, 11-16-06

Accomplished:

M. Software on all computers but Khan. A-O software on Data, Worf, Mr. Spock, Q, and Sarek. USB on these computers, too.

Khan, returned old hard drive and nuked. Q. needs a new network card. Worf has problem battery. Just need to hit esc. to start. Mr. Spock is having troubles starting; run Lucifer.
Once we started downloading the software, we realized that only Multplication master could be put on the computers. Computers wouldn’t let us download things, trial runs suddenly showed up that weren’t there when we downloaded the software, or the license agreement showed we couldn’t do it.

Monday (11-20-06) and Tuesday (11-21-06):

1. Finish A-O Software.
2. Finish USB software
3. Run Lucifer on Mr. Spock. Figure out what’s wrong.
4. New network card on Q.
5. Make sure all computers have anti-virus software.
6. Sound drivers

Monday, 11-20-06

Accomplished:

Khan is working; Spock is, too. Kahn has everything but USB and sound. Tried installing USB on Kahn; didn’t work. We think we might still have a bad A drive. Internet works on Khan, Data, Picard, and McCoy. Not working on Q; there may be a problem with the network card.
There are several steps to getting past the blasted MSN Internet set-up. Since IE is the default browser, we need to get past it. It took Martin awhile to figure it out.
We do have a problem: Geordi wouldn’t power on; no idea why.
Still no dice on sound card.

Tuesday (11-21-06):

Finish software and running USB on all computers. Check internet on all computers. Get Geordi working again. Anything else that needs doing.

Tuesday, 11-21-06

Accomplished:

See updated table for accomplishments. Got internet working on remaining computers (except Q and Spock). Switched out the power supply on Geordi; he’s now working again. Mr. Spock is having problems again. Since the burn in didn’t show any problems with the CPU, mother board, hard drive or memory, Martin suggested we wipe the hard drive and start over. We formatted and fdisked Mr. Spock but are having problems getting Windows 98 on the computer. We’ll have to fix this next week.

Next Week:

Finish.

Wednesday, 11-29-06

Accomplished:

The new diagram. (We had a problem because we’d orginally planned for two switches and got one hub instead. We have to calculate new cable lengths.)

Wednesday, 11-30-06

Accomplished:

Got Mr. Spock and Q. working. (Bobbi, what exactly did you do on Q and Spock to get them working? I’d like to know exactly what you did today so I can have an accurate record of it.)
We made nine cables so those should be finished (green/white, green, orange/white, blue, blue/white, orange, brown/white, brown). Also, someone in the class recommendedhttp://www.wartoft.nu/software/ for free software. We’ll check it out and see if we can find some more software for our comptuers.
We updated the diagram. Here’s the new and final version.

Next Week:

1. Make sure the hub works.
2. Make sure keyboards, mice work.
3. Collect power supplies; work?
4. Sound?
5. More software?
6. Anything else?
7. Transfer information to the wiki for Martin.

Wednesday, 12-6-06

See check lists.

We don’t have enough power cords (15; we’re supposed to have 19). We also don’t know where the monitors are. We e-mailed Martin. Mr. Spock is still freezing during start-up; too late now to fix him. (We did what we could so no need to get upset over it.) We can make a note of it on the computer. Cleaned mice and keyboards. (They were disgusting.)

Tomorrow:

1. Print out check lists.
2. Make two more labels
3. Two cars?
4. Get sleep.

Thursday, 12-7-06

Installation was a success.